OpenStudy (anonymous):

Snoopy is in his plane 4500 feet above the ground. He fires a projectile straight up. The initial velocity of the projectile is 340 feet per second. (After the projectile is fired, it has no acceleration, so its height can be modeled by a "falling body" model.) The Red Baron is maintaining a constant altitude of 5200 feet. When, on the way up, might the projectile intersect the Red Baron? Explain how you determined this. Use the eqution y=4500+340x-16x^2

OpenStudy (anonymous):

do you know how to find the vertex of the function?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

have you learned completing the square? using completing the square you can change the function into the vertex form \(y=a(x-h)^2-k\) where (h,k) is the vertex. the value of \(k\) will be the maximum height of the projectile. comparing it to the height of the Red Baron will determine whether it will intersect the projectile...

OpenStudy (anonymous):

No, we haven't gotten to that yet but I'll ask my teacher about that. Thank you!

OpenStudy (jdoe0001):

or you could also get the vertex of a parabola like so \(\bf y=ax^2+bx+c\) at the coordinates of \(\bf \left(-\cfrac{b}{2a}\quad ,\quad c-\cfrac{b^2}{4a}\right)\)

OpenStudy (anonymous):

Thank you!

OpenStudy (jdoe0001):

yw